A galvanometer of resistor 10 ohms gives a full scale deflection when a current of 50amphere flow through it.


How can it be converted to an ammeter to measure a current of 5Amphere?


How can it be converted to a voltmeter to measure a voltage of 3volts?

1
Expert's answer
2021-11-18T12:21:55-0500

For the ammeter the parallel shunt resistance is:


"R_S=\frac{I_g}{I-I_g}R_g=\frac{5}{50-5}10=1.11\Omega."

For the voltmeter the series resistance to be added is


"R_S=R_g\bigg(\frac{V}{V_V}-1\bigg)=10\bigg(\frac{10·50}{3}-1\bigg)=1657\Omega."